Part of an ANOVA table is shown below.

Source of
Variation

Sum of
Squares

Degrees of
Freedom

Mean
Square

F

Between Treatments

180

3

Within Treatments (Error)
TOTAL

480

18

The mean square due to error (MSE) is

a.

60.

b.

15.

c.

20.

d.

18.
Answer #1

Error Sum of squares = 480 - 180 = 300

Error degrees of freedom = 18 - 3 = 15

Hence,

Mean square due to error = 300/15 = 20

Option C is correct.